# Large Language Model Unlearning via Embedding-Corrupted Prompts

### v1

@@ 4.1 Prompt Classifier
For each unlearning task, we fine-tune a RoBERTa [79] to use as the prompt classifier on the corresponding $D_{r}$ and $D_{f}$. In entity and copyrighted content unlearning tasks, we use the entire $D_{f}$ to train the classifier [^4] In [87], the entire forget set is used during unlearning. because the unlearning target is fully captured by the forget set, which does not require generalization outside the set. For WMDP and MMLU, we only use a surrogate synthetic forget set $D_{\bar{f}}$ to train the prompt classifier, and the actual forget set $D_{f}$ is not accessible until evaluation. For all prompt classifiers, we use an independent validation set $D_{\textrm{val}}$ to tune the decision threshold $\tau$ and hyperparmaeters or to calibrate the empirical quantile $\hat{q}$, which is used to determine conformity. In Tables 6, 7 and 5, we show that all classifiers can distinguish $D_{r}$ and $D_{f}$ well, and generalize to unseen $D_{g}$ with low false positive rate. We provide further detailed information on how prompt classifiers are trained for each of the task and their performance in Section C.3.
Figure 2: Model utility versus forget quality (p-value) on three different forget set sizes of the TOFU dataset after unlearning. We show two models, Phi-1.5 (top) and Llama-2-7B-Chat (bottom). For GA, GD, KL, PO, and the prompting baseline, the forget quality are either too small or comes at the cost of substantial decrease of model utility. Negative preference optimization (NPO) [142] variants achieve a great balance in some cases, but the trade-off on model utility is still non-trivial. ECO-RN (random noise) and ECO-ZO (zero-out) achieve almost identical distribution to the retained model while having no sacrifice in model utility.
@@ 4.2 Entity Unlearning
Experimental setup The TOFU dataset [87] is a synthetic question-answering dataset of author biographies. The goal is for an LLM trained on the complete dataset (all authors) to unlearn a fraction of fictitious authors (1/5/10%) while retaining knowledge about both 1) the remaining fictitious authors and 2) the real-world. To assess forgetting and retaining, we use two metrics proposed alongside the TOFU dataset, forget quality and model utility. Forget quality is represented by a p-value from a Kolmogorov-Smirnov (KS) test, where high value indicates high similarity in distribution between the output of the unlearned model and that of the retained model. Model utility assesses the model’s performance on the retained set and real-world knowledge. For a detailed description of all the metrics, refer to Section C.1.1. We conduct experiments with two corruption functions, random noise (RN) and zero-out (ZO). We include all baselines in [87], a prompting baseline, and the recently proposed negative preference optimization (NPO) [142]. We provide formulations of all baselines Section C.4.
@@
Baselines struggle to forget or collapse in utility. We also observe that GA, GD, KL, PO, and the prompting baseline exhibit minimal forgetting when the forget set size is small (i.e., 1%). Meanwhile, some of them have a substantial decrease and even a collapse in utility when the forget set is larger (5% and 10%). Methods based on negative preference optimization [142] demonstrate a noticeably stronger trade-off compared to other baselines, especially with the NPO-RT. Nevertheless, the effectiveness of the NPO variants varies across different models and forget set sizes, and the loss in model utility is non-trivial. We present the full results on all metrics and baselines in Table 13 and Table 14 in Section E.2.
@@ 4.3 Hazardous Knowledge Unlearning
Experimental setup For both WMDP [70] and MMLU subset unlearning tasks [52], we directly unlearn on pre-trained models. WMDP benchmark focuses on unlearning knowledge in biology, chemistry, and cybersecurity. In MMLU subsets unlearning, the goal is to unlearn three subjects and retain their closely related counterparts: economics (econometrics), law (jurisprudence), and physics (math), all requiring high-precision forgetting to resolve knowledge entanglement. In line with [70], we assess all models based on their multiple-choice accuracy. A successfully unlearned model should exhibit an accuracy near random guessing (25% for four-option multiple-choice questions). We employ the ECO-RN variant (random noise) as the corruption function for both tasks. We only optimize corruption strength $\sigma$ for Zephyr-7B on a set of 100 synthetic questions and answers generated by GPT-4 to ensure the real questions is not exposed during unlearning. We use the same corruption parameter $\sigma$ for all other models. We compare our method against LLMU [136], SCRUB [66], SSD [39], RMU [70], and a prompting baseline that instructs the model not to answer questions within the domain correctly.
@@
ECO unlearns at high-precision. On MMLU subsets unlearning, both ECO and RMU successfully unlearn the three chosen subjects (Table 2). However, RMU’s accuracy in econometrics and jurisprudence significantly decreases. This implies that RMU might be sensitive to the entanglement of knowledge in closely related subjects. In contrast, this entanglement poses no problem for ECO’s prompt classifier due to its low false positive rate in retain domain.
@@ 4.4 Copyrighted Content Unlearning
Experimental setup. We select Harry Potter and the Sorcerer’s Stone [106] and BBC News articles [^5] https://huggingface.co/datasets/RealTimeData/bbc_news_alltime [72] as the copyrighted content material for unlearning and unlearn models fine-tuned on the text corpus. For this task, our goal is to prevent the unlearned model from generating passage with high similarity to the original text. For both datasets, we verify that the models used cannot generated the original passage and the generated text has low similarity to the original passage. We first fine-tune a pre-trained model on the corresponding corpus and use it as the model subject to unlearning, and use the original pre-trained checkpoint as the retained model. We use the original passage as the reference text and measure the text similarity between the reference and the text generated by the unlearned model using four text similarity metrics outlined in Section C.1.3, which we denote as the average similarity gap (ASG). Following [136], we also compute the perplexity and unique toke ratio to assess the if the generated text remains meaningful and diverse. We compare our method to baselines in [87], SCRUB [66], and LLMU [136]. We present full experimental details in Appendix C.
@@
ECO maintains high similarity to the retained model. In Table 3, ECO achieves scores sufficiently close to those of the retained model in terms of generated text similarity. On the general utility metric, our prompt classifiers effectively distinguish copyrighted content from general domain queries with no performance loss. KL minimization and LLMU are strong baselines in terms of similarity gap and general utility, but diversity in the generated text decreases after unlearning. Both gradient difference and random mismatch reduce the issue of model collapse but still lead to notable performance losses in general utility.
We further validate our findings on a total of 19 models in Section E.5 from Table 19 to Table 56. We observe that some baselines cannot consistently maintain strong results in either unlearning or general utility, while ECO remains stable and consistently achieves a low similarity gap with the retained model and unharmed utility.
@@ 5 Related Work
Unlearning for LLMs. Most existing machine unlearning methods for LLMs follow the traditional machine unlearning approaches [18, 96, 113] to minimize the influence of the forget samples via gradient updates. The most straightforward approach employ a mixture of forgetting and retaining objectives by performing gradient ascent updates on the non-desirable sequences and regular gradient descent on the desirable sequences [130, 136, 22, 135, 70, 142, 59]. Other methods identify and modify a small fraction of the weights responsible for the undesired behavior [132, 13, 55], or via weight arithmetic [141, 53, 97, 80]. The above optimization-based methods all require compute that scale with the model size. Our method leaves the LLM subject to unlearning intact and unlearns via steering the inputs to match the output distribution of a retained model. Compute-wise, our unlearning method is independent of the model size.
LLM guardrails. Guardrailing, which accesses prompts before using them as inputs to the model, has been widely applied on modern LLMs to prevent adversaries with harmful incentives [105, 56, 138, 88, 68, 31, 131, 56, 44, 24]. Our work is most related to in-context unlearning [101] and a recent guardrail baseline via prompting [122], both require no additional fine-tuning to achieve unlearning to some extent. [101] leverages modern LLM’s ability in in-context learning by prepending a small number of positive and negative samples in the prompt to steer the model’s response based on those samples. [122] guards the unlearning target via prompt injection, which inserts fixed instructions in the prompt to the LLM. Both method can only be applied to instruction-tuned models and rely on an LLM’s ability to follow instructions. Prepending such instructions also leads to significant performance degradation on regular tasks, as shown in [122].
Jailbreak via adversarial embeddings. Prior work on LLM jailbreaking [148, 38, 73, 41, 100] have demonstrated the power of adversarially optimizing toward a prompt that elicit a desired LLM response. In particular, [38] shows that the attack can be simplified to learning perturbation vectors added to the token embeddings, which eliminates the need to optimize over discrete tokens. Our results on the behavior of the attacked models are similar to the findings in [41], where inserting certain non-natural language token sequences in the prompt could elicit refusal behavior or incorrect answers from an instruction-tuned LLM. While jailbreak approaches can in theory be applied in unlearning applications, they are prohibitorily expensive to run during [78] and the additional requirement for specifying a sequence of desirable tokens. Both requirements make them unsuitable for the task of unlearning.
@@ 6 Conclusion
In this paper, we presented Embedding-COrrupted (ECO) Prompts, a novel method designed to address the dual challenges of knowledge entanglement and unlearning efficiency in LLM unlearning. ECO uses a thresholded prompt classifier to explicitly identify if a prompt is within the scope of the unlearning target. By decoupling the unlearning procedure from the LLMs themselves, ECO provide a scalable and efficient solution for unlearning in LLMs, paving the way for responsible and safe AI deployment in real-world applications.
Limitations and Broader Impacts We discuss the broader impacts and limitations of our approach in detail in Appendix A and Appendix B, respectively.
@@ Acknowledgement
We are thankful for the computing resources provided by the Pacific Research Platform’s Nautilus cluster, supported in part by National Science Foundation (NSF) awards CNS-1730158, ACI-1540112, ACI-1541349, OAC-1826967, OAC-2112167, CNS-2100237, CNS-2120019, the University of California Office of the President, and the University of California San Diego’s California Institute for Telecommunications and Information Technology/Qualcomm Institute, and CENIC for the 100Gbps networks.
@@ C.1.1 TOFU
The truth ratio is computed as the geometric mean [^6] We veified with the authors of the TOFU paper that the numerator should be a geometric mean instead of arithmetic mean, even though their paper still shows the arithmetic mean, at the time of writing this paper. This can also be confirmed by their implementation, which uses geometric mean: https://github.com/locuslab/tofu/blob/36811054f2376560c8d6629667059f3000e5603c/evaluate_util.py#L59. of multiple perturbed (incorrect) answers’ ($\mathcal{A}=\{\tilde{a}_{1},\tilde{a}_{2},...\}$) probabilities over the normalized conditional probability of the paraphrased answer $\hat{a}$.
$R_{\text{truth}}=\frac{\left(\prod_{i=1}^{|\mathcal{A}|}P(\tilde{a}\mid q)^{|1/\tilde{a}_{i}|}\right)^{1/|\mathcal{A}|}}{P(\hat{a}\mid q)^{1/|\hat{a}|}}$
@@ C.1.2 WMDP and MMLU
For both the WDMP [70] and MMLU subsets [52] unlearning, we employ multiple-choice accuracy as the primary evaluation metric. The underlying assumption is that a model unlearned on the target subject should demonstrate random-guessing accuracy on the task. Since both WMDP and MMLU consist of four-option multiple-choice questions, an accuracy close to 0.25 indicates successful unlearning. For each question, we adhere to the approach outlined by [70] and utilize the template provided in LABEL:lst:wmdp_template in a zero-shot manner. To derive the answer predicted by the LLM, we extract the logit scores corresponding to tokens [A, B, C, D] from the logit of the last token in the input sequence. The option with the highest logit score is deemed as the predicted answer.
@@
We also incorporate a probing evaluation, as done in [70], which trains a four-way linear probe on half of the data points from the biology, chemistry, and cybersecurity subsets. Specifically, a linear classifier is trained to predict the correct answer of the multiple-choice question based on the unlearned model’s output logits. The trained linear probe is then used to make predictions on the other half of the data points to infer the correct labels given the output logits. Successful unlearning should result in random-chance accuracy for the linear probe, which is 0.25 in our case.
@@ C.1.3 Harry Potter Book and BBC News Articles
We employ four text similarity metrics outlined below. For each metric, we use the original text (from the copyrighted material) as the reference and compute the similarity between the reference and the text generated by the LLM. A retained model that has never been trained on the reference text should have low similarity scores on all metrics, and a successfully unlearned model should have scores similar to that of the retained model. For both datasets, we evaluate similarity based on the first 256 tokens generated. This aligns with our fine-tuning setup in Section C.2.
@@
Following [136], we use the perplexity score and the unique token ratio measured on the generated text to assess fluency and diversity of the generated text. The perplexity is calculated by a reference model that has been fine-tuned on the target copyrighted content material. A sufficiently low perplexity indicates that the generated text might still be meaningful. The unique token is calculated using the tokenizer

### v2

@@ 4.1 Prompt Classifier
For each unlearning task, we fine-tune a RoBERTa [82] or a Llama-3.1-1B-Instruct [34] as the prompt classifier on the corresponding $D_{r}$ and $D_{f}$. In entity and copyrighted content unlearning tasks, we use the entire $D_{f}$ to train the classifier [^3] In [90], the entire forget set is used during unlearning. because the unlearning target is fully captured by the forget set, which does not require generalization outside the set. For WMDP and MMLU, we only use a surrogate synthetic forget set $D_{\bar{f}}$ to train the prompt classifier, and the actual forget set $D_{f}$ is not accessible until evaluation. For all prompt classifiers, we use an independent validation set $D_{\textrm{val}}$ to tune the decision threshold $\tau$ and hyperparameters or to calibrate the empirical quantile $\hat{q}$, which is used to determine conformity. In Tables 6, 7 and 5, we show that all classifiers can distinguish $D_{r}$ and $D_{f}$ well, and generalize to unseen $D_{g}$ with a low false positive rate. Meanwhile, although we do not specifically target out-of-distribution prompts related to jailbreak attempts, we demonstrate our classifiers’ ability to prevent such risks in section C.3.3. We also show that simple data augmentation techniques can further enhance our classifiers’ performance in detecting out-of-distribution and jailbreak prompts. We provide further detailed information on how prompt classifiers are trained for each task and their performance in Section C.3.
Figure 2: Model utility versus forget quality (p-value) on three different forget set sizes of the TOFU dataset after unlearning. We show two models, Phi-1.5 (top) and Llama-2-7B-Chat (bottom). For GA, GD, KL, PO, and the prompting baseline, the forget qualities are either too small or come at the cost of a substantial decrease in model utility. Negative preference optimization (NPO) [149] variants achieve a good balance in some cases, but the trade-off in model utility is still non-trivial. ECO-RN (random noise) and ECO-ZO (zero-out) achieve an almost identical distribution to the retained model while incurring no sacrifice in model utility.
@@ 4.2 Entity Unlearning
Experimental setup The TOFU dataset [90] is a synthetic question-answering dataset of author biographies. The goal is for an LLM trained on the complete dataset (all authors) to unlearn a fraction of fictitious authors (1/5/10%) while retaining knowledge about both 1) the remaining fictitious authors and 2) the real world. To assess forgetting and retention, we use two metrics proposed alongside the TOFU dataset: forget quality and model utility. Forget quality is represented by a p-value from a Kolmogorov-Smirnov (KS) test, where a high value indicates high similarity in distribution between the output of the unlearned model and that of the retained model. Model utility assesses the model’s performance on the retained set and real-world knowledge. For a detailed description of all the metrics, refer to Section C.1.1. We conduct experiments with two corruption functions, random noise (RN) and zero-out (ZO). We include all baselines from [90], a prompting baseline, and the recently proposed negative preference optimization (NPO) [149]. We provide formulations of all baselines in Section C.4.
@@
Baselines struggle to forget or collapse in utility. We also observe that GA, GD, KL, PO, and the prompting baseline exhibit minimal forgetting when the forget set size is small (i.e., 1%). Meanwhile, some of them experience a substantial decrease and even a collapse in utility when the forget set is larger (5% and 10%). Methods based on negative preference optimization [149] demonstrate a noticeably stronger trade-off compared to other baselines, especially with NPO-RT. Nevertheless, the effectiveness of the NPO variants varies across different models and forget set sizes, and the loss in model utility is non-trivial. We present the full results on all metrics and baselines in Table 17 and Table 18 in Section E.2.
@@ 4.3 Hazardous Knowledge Unlearning
Experimental setup For both WMDP [73] and MMLU subset unlearning tasks [55], we directly unlearn on pre-trained models. The WMDP benchmark focuses on unlearning knowledge in biology, chemistry, and cybersecurity. In MMLU subset unlearning, the goal is to unlearn three subjects while retaining their closely related counterparts: economics (econometrics), law (jurisprudence), and physics (math), all requiring high-precision forgetting to resolve knowledge entanglement. In line with [73], we assess all models based on their multiple-choice accuracy. A successfully unlearned model should exhibit an accuracy near random guessing (25% for four-option multiple-choice questions). We employ the ECO-RN variant (random noise) as the corruption function for both tasks. We optimize the corruption strength $\sigma$ only for Zephyr-7B on a set of 100 synthetic questions and answers generated by GPT-4 to ensure that real questions are not exposed during unlearning. The same corruption parameter $\sigma$ is used for all other models. We compare our method against LLMU [143], SCRUB [69], SSD [41], RMU [73], and a prompting baseline that instructs the model not to answer questions within the domain correctly.
@@
ECO unlearns at high precision. On MMLU subset unlearning, both ECO and RMU successfully unlearn the three chosen subjects (Table 2). However, RMU’s accuracy in econometrics and jurisprudence significantly decreases. This implies that RMU might be sensitive to the entanglement of knowledge in closely related subjects. In contrast, this entanglement poses no problem for ECO’s prompt classifier due to its low false positive rate in the retain domain.
@@ 4.4 Copyrighted Content Unlearning
Experimental setup We select Harry Potter and the Sorcerer’s Stone [^4] We purchased the Harry Potter and the Sorcerer’s Stone ebook and extracted the entire corpus to train our HP Book prompt classifier. [112] and BBC News articles [^5] https://huggingface.co/datasets/RealTimeData/bbc_news_alltime [75] as the copyrighted content material for unlearning and unlearn models fine-tuned on the text corpus. For this task, our goal is to prevent the unlearned model from generating passages with high similarity to the original text. For both datasets, we verify that the models used cannot generate the original passage and that the generated text has low similarity to the original passage. We first fine-tune a pre-trained model on the corresponding corpus and use it as the model subject to unlearning, with the original pre-trained checkpoint serving as the retained model. We use the original passage as the reference text and measure the text similarity between the reference and the text generated by the unlearned model using four text similarity metrics outlined in Section C.1.3, which we denote as the average similarity gap (ASG). Following [143], we also compute the perplexity and unique token ratio to assess whether the generated text remains meaningful and diverse. We compare our method to baselines in [90], SCRUB [69], and LLMU [143]. We present full experimental details in Appendix C.
@@
ECO maintains high similarity to the retained model. In Table 3, ECO achieves scores sufficiently close to those of the retained model in terms of generated text similarity. On the general utility metric, our prompt classifiers effectively distinguish copyrighted content from general domain queries with no performance loss. KL minimization and LLMU are strong baselines in terms of similarity gap and general utility, but the diversity of the generated text decreases after unlearning. Both gradient difference and random mismatch reduce the issue of model collapse but still lead to notable performance losses in general utility.
We further validate our findings on a total of 19 models in Section E.5, spanning from Table 23 to Table 60. We observe that some baselines cannot consistently maintain strong results in either unlearning or general utility, while ECO remains stable and consistently achieves a low similarity gap with the retained model and unharmed utility.
@@ 5 Related Work
Unlearning for LLMs Most existing machine unlearning methods for LLMs follow traditional machine unlearning approaches [19, 101, 119] to minimize the influence of the forget samples via gradient updates. The most straightforward approach employs a mixture of forgetting and retaining objectives by performing gradient ascent updates on the non-desirable sequences and regular gradient descent on the desirable sequences [137, 143, 23, 142, 73, 149, 62]. Other methods identify and modify a small fraction of the weights responsible for the undesired behavior [139, 13, 58], or use weight arithmetic [148, 56, 102, 83]. The above optimization-based methods all require compute that scales with the model size. Our method leaves the LLM subject to unlearning intact and unlearns by steering the inputs to match the output distribution of a retained model. Compute-wise, our unlearning method is independent of the model size.
LLM guardrails Guardrailing, which accesses prompts before using them as inputs to the model, has been widely applied to modern LLMs to prevent adversaries with harmful incentives [111, 59, 145, 92, 71, 32, 138, 59, 47, 25]. Our work is most related to in-context unlearning [106] and a recent guardrail baseline via prompting [129], both of which require no additional fine-tuning to achieve unlearning to some extent. [106] leverages modern LLMs’ ability in in-context learning by prepending a small number of positive and negative samples in the prompt to steer the model’s response based on those samples. [129] guards the unlearning target via prompt injection, which inserts fixed instructions in the prompt to the LLM. Both methods can only be applied to instruction-tuned models and rely on an LLM’s ability to follow instructions. Prepending such instructions also leads to significant performance degradation on regular tasks, as shown in [129].
Jailbreak via adversarial embeddings Prior work on LLM jailbreaking [155, 40, 76, 44, 105] has demonstrated the power of adversarially optimizing toward a prompt that elicits a desired LLM response. In particular, [40] shows that the attack can be simplified to learning perturbation vectors added to the token embeddings, which eliminates the need to optimize over discrete tokens. Our results on the behavior of the attacked models are similar to the findings in [44], where inserting certain non-natural language token sequences in the prompt could elicit refusal behavior or incorrect answers from an instruction-tuned LLM. While jailbreak approaches can theoretically be applied in unlearning applications, they are prohibitively expensive to run [81], and there is an additional requirement for specifying a sequence of desirable tokens. Both requirements make them unsuitable for the task of unlearning.
@@ 6 Conclusion
In this paper, we introduced Embedding-COrrupted (ECO) Prompts, a novel method to tackle the dual challenges of knowledge entanglement and unlearning efficiency in LLMs. ECO leverages a thresholded prompt classifier to determine whether a prompt falls within the scope of the unlearning target. By decoupling the unlearning process from the LLMs themselves, ECO offers a scalable and efficient approach that remains effective across a wide range of model sizes, from 0.5B to 236B parameters, with minimal side effects and no additional computational overhead. Our experiments across three unlearning tasks validate ECO’s effectiveness, setting a foundation for responsible AI deployment in real-world scenarios.
@@ Acknowledgement
This work is partially supported by the National Science Foundation under grants IIS-2143895, IIS-2040800, IIS-2416896, IIS-2007951 and CCF-2023495. We are also thankful for the computing resources provided by the Pacific Research Platform’s Nautilus cluster, supported in part by National Science Foundation (NSF) awards CNS-1730158, ACI-1540112, ACI-1541349, OAC-1826967, OAC-2112167, CNS-2100237, CNS-2120019, the University of California Office of the President, the University of California San Diego’s California Institute for Telecommunications and Information Technology/Qualcomm Institute, and CENIC for the 100 Gbps networks.
@@ C.1.1 TOFU
The truth ratio is computed as the geometric mean [^6] We verified with the authors of the TOFU paper that the numerator should be a geometric mean instead of an arithmetic mean, even though their paper still shows the arithmetic mean at the time of writing this paper. This can also be confirmed by their implementation, which uses the geometric mean: https://github.com/locuslab/tofu/blob/36811054f2376560c8d6629667059f3000e5603c/evaluate_util.py#L59 of multiple perturbed (incorrect) answers’ ($\mathcal{A}=\{\tilde{a}_{1},\tilde{a}_{2},...\}$) probabilities over the normalized conditional probability of the paraphrased answer $\hat{a}$.
$R_{\text{truth}}=\frac{\left(\prod_{i=1}^{|\mathcal{A}|}P(\tilde{a}_{i}\mid q)^{1/|\tilde{a}_{i}|}\right)^{1/|\mathcal{A}|}}{P(\hat{a}\mid q)^{1/|\hat{a}|}}$
@@ C.1.2 WMDP and MMLU
For both the WDMP [73] and MMLU subsets [55] unlearning, we employ multiple-choice accuracy as the primary evaluation metric. The underlying assumption is that a model unlearned on the target subject should demonstrate random-guessing accuracy on the task. Since both WDMP and MMLU consist of four-option multiple-choice questions, an accuracy close to 0.25 indicates successful unlearning. For each question, we adhere to the approach outlined by [73] and utilize the template provided in LABEL:lst:wmdp_template in a zero-shot manner. To derive the answer predicted by the LLM, we extract the logit scores corresponding to tokens [A, B, C, D] from the logits of the last token in the input sequence. The option with the highest logit score is deemed the predicted answer.
@@
We also incorporate a probing evaluation, as done in [73], which trains a four-way linear probe on half of the data points from the biology, chemistry, and cybersecurity subsets. Specifically, a linear classifier is trained to predict the correct answer to the multiple-choice question based on the unlearned model’s output logits. The trained linear probe is then used to make predictions on the other half of the data points to infer the correct labels given the output logits. Successful unlearning should result in random-chance accuracy for the linear probe, which is 0.25 in our case.
@@ C.1.3 Harry Potter Book and BBC News Articles
We employ four text similarity metrics outlined below. For each metric, we use the original text (from the copyrighted material) as the reference and compute the similarity between the reference and the text generated by the LLM. A retained model that has never been trained on the reference text should have low similarity scores on all metrics, and a successfully unlearned model should have scores similar to those of the retained model. For both datasets, we evaluate similarity based on the first 256 tokens generated. This aligns with our fine-tuning setup in Section C.2.
@@
Following [143], we use the perplexity score and the unique token ratio measured on the generated text to assess the fluency and diversity of the generated text. The perplexity is calculated by a reference model that has been fine-tuned on the target copyrighted content material. A sufficiently low perplexity indicates that the generated text might still be meaningful. The unique token ratio is calculated as the number of unique token set over all tokens generated in the outputs.